Find the center and radius of the circle (x-5)2 + y2 + 2y=7: Find the center and radius of the circle (x-5)2 + y2 + 2y=7
Answer by Alan3354(1449)

Find the center and radius of the circle (x-5)2 + y2 + 2y=7
Put in the form (x-a)^2+(y-b)^2 = r^2
x is ok as it is.
Complete the square for y
y^2+2y+1 = (y+1)^2, so add 1 to both sides:
(x-5)^2+(y-(-1))^2 = 8
The center is at (a,b), or (5,-1) and the radius is sqrt(8).
sqrt(8) = 2sqrt(2)